# Look Around and Find Out: OOD Detection with Relative Angles

### v1

@@ 3.3 Features on the Decision Boundary
$\frac{d(\mathbf{z}-\bm{\mu_{\text{ID}}},\mathbf{z}_{db}-\bm{\mu_{\text{ID}}})}{\sin{(\theta)}}=\frac{d(\mathbf{z}-\bm{\mu_{\text{ID}}},0)}{\sin{(\alpha)}}$ (4)
$\frac{\sin{(\theta)}}{\sin{(\alpha)}}=\frac{d(\mathbf{z}-\bm{\mu_{\text{ID}}},\mathbf{z}_{db}-\bm{\mu_{\text{ID}}})}{d(\mathbf{z}-\bm{\mu_{\text{ID}}},0)}=\text{fDBD}(\mathbf{z})$ (5)
Figure 2: Histogram of ID (CIFAR-10) and OOD (Texture) with respect to the sine of the angle of the triangle that looks at the edge $\mathbf{z}-\bm{\mu_{\text{ID}}}$. This empirically shows that $\sin{(\alpha)}$ is not highly informative to distinguish between ID or OOD.
where $\theta$ and $\alpha$ are the angles opposite to the sides $\mathbf{z}-\bm{\mu_{\text{ID}}}-\mathbf{(z_{db}}-\bm{\mu_{\text{ID}}})$ and $\mathbf{z}-\bm{\mu_{\text{ID}}}$ respectively. Although the observation they made on comparing the distances to the decision boundaries at equal deviation levels from the mean of in-distribution is inspiring, we claim that the angle $\alpha$ is not very informative for ID and OOD separation. This is because $\alpha$ is connected to the magnitude of the feature vector relative to $\bm{\mu_{\text{ID}}}$, which may not directly correlate with OOD characteristics. On Figure 2 we show the $\sin{(\alpha)}$ values between CIFAR-10  [20] and Texture  [5] datasets, empirically justifying that including this term impedes fDBD’s performance. Omitting the denominator from Equation 5 allows to effectively capture the relation between the feature representation and the decision boundary from the mean of in-distribution’s view.
@@ 4 Experiments
In this section, we demonstrate the performance of LAFO on various settings and benchmarks. We will first provide our results on the most common benchmarks CIFAR-10  [20] and ImageNet  [6] to show small and large scale performance of our scoring. Then, we show the flexibility of LAFO by (i) combining it with different activation shaping algorithms, (ii) using it to aggregate different architectures’ confidences via simply summing their scores, exploiting the scale-invariance property of LAFO. Moreover, an ablation study is presented to show the efficacy of the design choices.
@@
Benchmarks: We mainly consider two widely used benchmarks: CIFAR-10  [20] and ImageNet  [6]). We included the evaluation on CIFAR-10 OOD Benchmark to show the performance on smaller scale datasets. In CIFAR-10 experiments, we use a pretrained ResNet-18 architecture  [14] trained with supervised contrastive loss [19], following previous literature [25, 39, 36]. During inference $10.000$ test samples are used to set the in-distribution scores and choose the threshold value $\lambda$; while the datasets SVHN  [33], iSUN  [48], Places365  [52] and Texture  [5] are used to obtain out-of-distribution scores and metric evaluation. Similarly, for large scale ImageNet OOD Benchmark, we use a pretrained ResNet-50 architecture trained with supervised contrastive loss. In-distribution validation set of size $50.000$ is used to set the ID scores and the threshold, and the datasets iNaturalist  [42], SUN  [45], Places365  [52] and Texture  [5] are used to obtain out-of-distribution scores and metric evaluation. We analyzed the case of architectures trained with Cross Entropy loss, in the ensemble experiment of Table 3. Note that, in this typical OOD Detection Benchmarks the samples that have same classes as ID are removed from their OOD counterparts, following the work  [17] and fitting into our problem setting.
@@
Table 1: LAFO achieves state-of-the-art performance on CIFAR-10 OOD benchmark. Evaluated on ResNet-18 with FPR95 and AUROC. $\uparrow$ indicates that larger values are better and vice versa. Best performance highlighted in bold. Methods with * are hyperparameter-free.
@@
LAFO * 3.53 99.16 8.36 98.28 23.40 94.88 8.58 98.34 10.97 97.67
4.1 OOD Detection on CIFAR-10 and ImageNet Benchmarks
Table 1 and Table 2 shows the performance of LAFO along with the 9 baselines on CIFAR-10 and ImageNet OOD Benchmarks, respectively. All the baselines on CIFAR-10 use ResNet-18 architecture and on ImageNet use ResNet-50. We compare with methods in three categories : (i) logit-based score functions, (ii) methods that utilize contrastively learned representations (iii) hyperparameter free methods. Our proposed method reaches state-of-the-art performance on both benchmarks, reducing the FPR95 on average by $7.74\%$ on Imagenet and $0.88\%$ on CIFAR10. In the following, we provide a detailed analysis of these results.
Table 2: LAFO achieves state-of-the-art performance on ImageNet OOD benchmark. Evaluated on ResNet-50 with FPR95 and AUROC. $\uparrow$ indicates that larger values are better and vice versa. Best performance highlighted in bold. Methods with * are hyperparameter-free.
@@ 4.1 OOD Detection on CIFAR-10 and ImageNet Benchmarks
LAFO * 12.27 97.42 31.80 92.85 40.71 90.10 15.39 96.68 25.04 94.26
@@
LAFO continues to show the success of distance-based methods over logit-based methods. Logit-based scoring methods MSP  [15], Energy  [26] are one of the earliest baselines proving their success on measuring model’s confidences. MSP measures the maximum softmax probability as its score while Energy does a logsumexp operation on the logits. Recent distance-based methods like KNN+  [39] and fDBD  [25] outperforms the early logit-based ones. Similarly, LAFO achieves significantly better performance on both benchmarks, reducing the FPR95 up to $49.81\%$ and $41.91\%$ while improving the AUROC up to $7.53\%$ and $12.27\%$ on CIFAR-10 and ImageNet OOD benchmarks.
@@
LAFO improves on the recent success of methods using contrastively learned features. Table 1 and 2 show the success of recent methods CSI  [40], SSD+  [36], KNN+  [39] and fDBD  [25] that utilizes contrastively learned representations over the ones those do not use. We observe that the additional structure the supervised contrastive loss puts on the feature representations are particularly beneficial to the distance-based methods. LAFO also benefits from more structured representations on the feature space, as it explores the relationship between the representation and the decision boundaries. Notably, LAFO improves both of the metrics on both CIFAR-10 and ImageNet benchmarks, achieving the state-of-the-art performance.
LAFO enjoys being hyperparameter-free while offering state-of-the-art performance. Methods that use hyperparameters require a holdout set the tune them. Moreover, having different optimal hyperparameters for different benchmarks makes it harder to use them in real world applications. MSP  [15], Energy  [26] and fDBD  [25] are the baselines which do not require any hyperparameters. LAFO outperforms the most competitive hyperparameter-free baseline fDBD by having $10.97\%$ FPR95 on CIFAR-10 as opposed to fDBD’s $11.85\%$, and $25.04\%$ FPR95 on ImageNet compared to fDBD’s $32.78\%$.
4.2 Model Ensembling with LAFO
Recent works  [49] and  [46] show that creating an ensemble of models can enhance the OOD performance. Inspired from these works, and from the observation that scale invariant representations are compatible between distinct models [32], we demonstrate that scale-invariant score functions can aggregate the confidences from different models, by simply summing their scores. On Table 3 we show the individual performances of models ResNet-50, ResNet-50 with supervised contrastive loss and ViT-B/16 as well as their combined performances using the scale-invariant LAFO.
Figure 3: LAFO can be used for ensemble OOD detection due to its scale-invariance property. Evaluated on ImageNet OOD benchmark. Best performance highlighted in bold.
Method Avg
[…]
fDBD w/ResNet50 51.35 89.20
fDBD w/ResNet50-supcon 32.78 92.86
fDBD w/ViT-B/16 41.55 91.05
LAFO w/ResNet50 44.58 90.68
LAFO w/ResNet50-supcon 25.04 94.26
LAFO w/ViT-B/16 39.92 91.38
Ensemble fDBD 31.05 95.29
Ensemble LAFO 22.53 96.41
Note that, we show the scale-invariance property of fDBD on section 3.3 and added it to demonstrate the same idea and also to compare with LAFO. It can be seen that for both of the score functions, the performance of ensemble is better than their individual counterparts showing that score aggregation improves their OOD performance. Moreover, the ensemble with LAFO achieves a performance with $22.53\%$ FPR95 and $96.41\%$ AUROC, improving the metrics compared to the best individual performer in the ensemble by $2.51\%$ and $2.15\%$ respectively. We show qualititave evidence of the improved performance by plotting the ID and OOD histograms on ImageNet  [6](ID) and iNaturalist  [42] (OOD) datasets, respectively, in Figure 6 in the Appendix, demonstrating a better separation in the ensembled model. In summary, we demonstrate that scale-invariance of LAFO allows aggregating different models’ confidences to solve OOD Detection Problem.
Table 3: LAFO can be used as a plug-in on top of activation shaping algorithms. Evaluated under ImageNet OOD benchmark. $\uparrow$ indicates that larger values are better and vice versa. Best performance highlighted in bold.

### v2

@@ 3.2 Features on the Decision Boundary
$\frac{d(\mathbf{z}-\bm{\mu_{\text{ID}}},\mathbf{z}_{db}-\bm{\mu_{\text{ID}}})}{\sin{(\theta)}}=\frac{d(\mathbf{z}-\bm{\mu_{\text{ID}}},0)}{\sin{(\alpha)}}\Rightarrow\frac{\sin{(\theta)}}{\sin{(\alpha)}}=\frac{d(\mathbf{z}-\bm{\mu_{\text{ID}}},\mathbf{z}_{db}-\bm{\mu_{\text{ID}}})}{d(\mathbf{z}-\bm{\mu_{\text{ID}}},0)}=\text{fDBD}(\mathbf{z})$ (4)
where $\theta$ and $\alpha$ are the angles opposite to the sides $\mathbf{z}-\bm{\mu_{\text{ID}}}-\mathbf{(z_{db}}-\bm{\mu_{\text{ID}}})$ and $\mathbf{z}-\bm{\mu_{\text{ID}}}$ respectively. Although the observation they made on comparing the distances to the decision boundaries at equal deviation levels from the mean of in-distribution is inspiring, we claim that the angle $\alpha$ is not very informative for ID and OOD separation. This is because $\alpha$ is connected to the magnitude of the feature vector relative to $\bm{\mu_{\text{ID}}}$, which may not directly correlate with OOD characteristics. On Figure 2 we show the $\sin{(\alpha)}$ values between CIFAR-10  (Krizhevsky, 2009) and Texture  (Cimpoi et al., 2014) datasets, empirically justifying that including this term impedes fDBD’s performance. Omitting the denominator from Equation 4 allows to effectively capture the relation between the feature representation and the decision boundary from the mean of in-distribution’s view. See Appendix  J for more detailed explanation on the performance relationship between raw distance, fDBD and ORA.
Figure 2: Histogram of ID (CIFAR-10) and OOD (Texture) samples with respect to the sine of the angle formed with the vector $\mathbf{z}-\bm{\mu}_{\text{ID}}$. This empirically shows that $\sin{(\alpha)}$ is not highly informative for distinguishing ID from OOD.
@@ 4 Experiments
In this section, we first test ORA on a large-sale ImageNet OOD benchmark (Deng et al., 2009) that spans nine models –including ConvNeXt (Liu et al., 2022), Swin (Liu et al., 2021), DeiT (Touvron et al., 2021) and EVA (Fang et al., 2023)– to establish the performance beyond the usual ResNet-50 (He et al., 2016) setting. We then demonstrate how ORA benefits from contrastively learned features (i) on CLIP (Radford et al., 2021) in both zero-shot and linear-probe modes, and (ii) on CIFAR-10 (Krizhevsky, 2009) and ImageNet with ResNet18/50 checkpoints trained with supervised contrastive (SCL) loss (Khosla et al., 2020). Next, we show that ORA’s scale-invariant scores can be ensemble-summed across architectures for additional gains, and that it pairs seamlessly with post-hoc activation-shaping methods such as ReAct and ASH. Finally, we present an ablation study to asses the contributions of key design choices.
Benchmarks.
@@
We consider two widely used benchmarks: CIFAR-10  (Krizhevsky, 2009) and ImageNet  (Deng et al., 2009)). We included the evaluation on CIFAR-10 OOD Benchmark to show the performance on smaller scale datasets. In CIFAR-10 experiments, we use a pretrained ResNet-18 architecture  He et al. (2016) trained with supervised contrastive loss (Khosla et al., 2020), following previous literature Liu and Qin (2024); Sun et al. (2022); Sehwag et al. (2021). During inference $10.000$ test samples are used to set the in-distribution scores and choose the threshold value $\lambda$; while the datasets SVHN  (Netzer et al., 2011), iSUN  (Xu et al., 2015), Places365  (Zhou et al., 2017) and Texture  (Cimpoi et al., 2014) are used to obtain out-of-distribution scores and metric evaluation.
For the large-scale ImageNet OOD benchmark, we extend prior evaluations  (Liu and Qin, 2024; Sun et al., 2022; Park et al., 2023; Ren et al., 2021; Sun et al., 2021; Xu et al., 2024b) by going beyond ResNet-50 and including a diverse set of nine models, such as ConvNeXt (Liu et al., 2022), Swin Transformer (Liu et al., 2021), DeiT (Touvron et al., 2021), and EVA (Fang et al., 2023), along with their ImageNet-21k pretrained counterparts when available. This broader evaluation allows a more comprehensive assessment of OOD detection performance across modern architectures. A validation set of 50,000 ImageNet samples is used to set ID scores and the threshold, while the OOD datasets include iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), Places365 (Zhou et al., 2017), and Texture (Cimpoi et al., 2014). Note that, in this typical OOD Detection Benchmarks the samples that have same classes as ID are removed from their OOD counterparts, following the work  Huang and Li (2021) and fitting into our problem setting.
Metrics.
We report two metrics in our experiments: False-positive rate at $\%95$ true positive rate (FPR95), and Area Under the Receiver Operating Characteristic Curve (AUROC). FPR95 measures what percentage of OOD data we falsely classify as ID where our threshold includes $95\%$ of ID data. Therefore, smaller FPR95 indicates a better performance by sharply controlling the false positive rate. On the other hand, AUROC measures the model’s ability to distinguish between ID and OOD by calculating the area under the curve that plots the true positive rate against the false positive rate across thresholds. AUROC shows how rapidly we include ID data while paying the cost of including false positives. Thus, higher AUROC shows a better result.
Table 1: Average FPR95 (%) on the ImageNet OOD detection benchmark. Each column corresponds to a different detection method, and each row to a model backbone. Lower values indicate better performance. The best result per row is highlighted in bold.
Model MSP MaxLogit Energy R-Mah R-Cos P-Cos NNG KNN+ fDBD NCI Maha+ NECO ORA ConvNeXt 63.54 60.69 78.70 45.79 46.99 49.37 39.54 47.40 39.03 36.44 36.68 34.52 34.92 ConvNeXt-pre 45.57 42.39 46.95 34.91 36.16 35.47 76.91 77.23 42.13 50.48 37.50 34.40 29.53 Swin 61.99 59.21 65.78 45.25 46.26 49.37 57.34 53.79 52.34 57.25 48.21 58.24 51.56 Swin-pre 46.20 40.86 42.33 39.40 35.66 33.94 44.06 51.34 38.92 39.63 42.67 36.49 35.13 DeiT 57.48 57.52 70.05 52.51 47.63 61.62 58.02 71.34 55.98 58.11 54.18 58.49 54.16 DeiT-pre 57.45 54.72 55.56 39.37 38.49 39.11 43.56 44.18 46.69 57.76 38.68 52.61 37.76 EVA 43.75 40.37 43.40 27.45 27.94 27.68 37.33 34.19 36.10 39.34 31.11 36.05 33.37 ResNet-50 66.95 64.29 58.41 83.29 59.51 45.53 55.22 53.97 51.19 63.03 49.67 56.62 43.63 ResNet-50 (SCL) 50.06 97.59 98.46 77.52 48.18 31.57 98.80 38.47 32.78 96.02 45.06 97.58 25.04 Average 54.78 57.52 62.18 49.50 42.98 41.52 56.75 52.43 43.90 55.34 42.64 51.67 38.34
4.1 OOD Detection on ImageNet Benchmark
Table 1 presents an extensive evaluation of OOD detection performance on the ImageNet benchmark, spanning 9 diverse model backbones and 12 scoring methods. In contrast to prior work that typically centers evaluation on ResNet-50Liu and Qin (2024); Sun et al. (2022); Park et al. (2023); Ren et al. (2021); Sun et al. (2021); Xu et al. (2024b), we expand the setting to include modern transformer-based architectures—ConvNeXt, Swin, DeiT, and EVA—as well as their ImageNet-21k pretrained variants when available. This large-scale comparison provides a better performance evaluation across architectures.
ORA consistently demonstrates strong performance across the board. It achieves the best average FPR95 across all models at $38.34\%$, outperforming the second-best method by a margin of $3.18\%$ and the best FPR95 in 5 out of 9 model settings. Notably, ORA achieves the single best FPR95 score overall with $25.04\%$ on the ResNet-50 (SCL) model, outperforming the next-best score of $27.45\%$ obtained by R-Mah with EVA by a margin of $2.41\%$.
Furthermore, in the few cases where ORA is not the top-performing method—such as with vanilla Swin or DeiT models—there is no alternative method that clearly dominates. The best-performing method varies across these settings (R-Mah, R-Cos and P-Cos), indicating that no other baseline consistently adapts well to different backbones. In contrast, ORA remains within the top three methods across all model architectures, demonstrating both robustness and versatility. These results highlight ORA’s ability to generalize across both convolutional and transformer-based models, pretrained or otherwise, and establish it as a strong state-of-the-art approach for OOD detection on ImageNet.
Near-OOD Results.
In Appendix I we present complementary near-OOD evaluations on the NINCO (Bitterwolf et al., 2023) and SSB-Hard (Vaze et al., 2022) datasets. Because these benchmarks feature higher class similarity between ID and OOD samples, all methods show degraded separability compared to far-OOD settings and performance differences are more subtle. In this regime, approaches that explicitly use class centroids tend to be more effective (albeit with higher memory and information requirements), especially when OOD examples are visually close to certain ID classes. Although ORA does not rely on individual class means, it achieves competitive performance on par with centroid-based methods. Consistent with Bitterwolf et al. (2023), we also observe that no single method emerges as clearly state-of-the-art in this benchmark.
4.2 OOD Detection with CLIP
In this section, we evaluate the performance of post-hoc OOD detection methods on the CLIP architecture. CLIP, with its pre-trained vision-language capabilities, provides a foundation for OOD detection. We explore two approaches: linear probing and a zero-shot extension that leverages CLIP’s inherent structure to avoid additional training.
Linear probing with CLIP.
For methods requiring decision boundaries (e.g. Hendrycks and Gimpel (2022); Liu et al. (2020); Liu and Qin (2024), and ORA), we used the vision encoder of CLIP to extract features and trained a linear probe for the classification task. While effective, this reduces flexibility due to the need for additional training.
@@ 4.2 OOD Detection with CLIP
[…]
Zero-shot OOD detection with CLIP.
To mitigate the limitation of requiring decision boundaries, we derive a zero-shot extension that defines them using text embeddings. For vision-language models like CLIP, we define the decision boundaries by vectors whose cosine similarity to the text embeddings of two classes are equal. Since cosine similarity is scale-invariant, these vectors are unique up to its norm.
Table 3: OOD detection performance on the ImageNet benchmark using CLIP-ViT-H/14 features. We report FPR95 (%) and AUROC (%) for four OOD datasets under two regimes: linear probing (lp) and zero-shot (zs). The best result per column is highlighted in bold.
Method iNaturalist SUN Places Texture Avg
[…]
MSP/lp 15.74 96.64 46.00 88.68 48.73 87.40 40.87 87.98 37.83 90.18
Energy/lp 7.26 97.94 34.62 92.13 41.32 90.05 37.02 90.98 30.06 92.77
MaxL/lp 9.02 97.91 37.15 91.70 42.25 89.86 36.37 90.80 31.20 92.57
NNG/lp 7.26 97.94 34.62 92.12 41.32 90.05 37.02 90.98 30.06 92.77
KNN/zs 80.20 87.86 86.68 84.63 73.51 86.07 70.27 84.60 77.66 85.79
fDBD/zs 9.31 98.11 22.32 94.78 29.15 93.20 40.12 90.25 25.23 94.08
fDBD/lp 5.62 98.48 32.18 93.89 35.74 92.54 27.13 93.71 25.17 94.66
Pcos/zs 19.86 96.09 28.67 93.58 36.09 91.84 45.21 88.97 32.46 92.62
MCM/zs 17.81 96.52 26.04 93.94 35.07 91.54 41.12 90.92 30.01 93.23
RMh/zs 3.64 98.85 30.71 93.12 34.72 92.18 21.21 93.84 22.57 94.49
ORA/zs 14.12 97.41 22.97 94.97 28.01 93.41 38.28 90.73 25.85 94.13
ORA/lp 6.66 98.16 30.35 94.43 33.79 93.20 24.95 94.34 23.94 95.03
[…]
Positive effect of contrastive features.
An important observation from our results is the positive effect of contrastive learning on ORA’s performance. Contrastive objectives—whether within a single modality, such as SCL, or across modalities, as in CLIP’s image-text training—impose a geometric structure on the representation space that separates semantic concepts. ORA benefits particularly from this structure as confirmed by the results in Table 2: we show that ORA delivers the lowest average FPR95 on both benchmarks, cutting the previous best score from $11.85\%$ to $10.97\%$ on CIFAR-10 (-0.88 pp) and from $32.78\%$ to $25.04\%$ on ImageNet (-7.74 pp), thereby outperforming every competing method with SCL-trained ResNet-18 and ResNet-50 backbones. Similarly, in Table 3, ORA shows strong results in both the zero-shot and linear probe settings of CLIP. These findings suggest that ORA is particularly well-suited to contrastively structured representations, hinting at a deeper connection between contrastive learning and OOD detection performance.
4.3 Model Ensembling with ORA
Table 4: ORA can be used for ensemble OOD detection due to its scale-invariance property. Evaluated on the ImageNet OOD benchmark. Best performance is highlighted in bold.
[…]
Recent works  Xue et al. (2024) and  Xu et al. (2024a) show that creating an ensemble of models can enhance the OOD performance. Inspired from these works, and from the observation that scale invariant representations are compatible between distinct models (Moschella et al., 2023), we demonstrate that scale-invariant score functions can aggregate the confidences from different models, by simply summing their scores. On Table 4 we show the individual performances of models ResNet-50, ResNet-50 with SCL and ViT-B/16 as well as their combined performances using the scale-invariant ORA.
Note that we demonstrate the scale-invariance property of fDBD in Equation 4 and include it in ensemble experiments to compare with ORA. It can be seen that for both of the score functions, the performance of ensemble is better than their individual counterparts showing that score aggregation improves their OOD performance. Moreover, the ensemble with ORA achieves a performance with $22.53\%$ FPR95 and $96.41\%$ AUROC, improving the metrics compared to the best individual performer in the ensemble by $2.51\%$ and $2.15\%$ respectively. In summary, we demonstrate that scale-invariance of ORA allows aggregating different models’ confidences to solve OOD Detection Problem.
4.4 ORA with Activation Shaping Algorithms
Recent methods ReAct  Sun et al. (2021), ASH  Djurisic et al. (2023) and Scale  Xu et al. (2024b) show their success to modify the feature representations to reduce model’s overconfident predictions. All three methods adopt a hyperparameter percentile to choose how to truncate and scale the feature representations using ID data statistics.
Table 5: ORA can be used as a plug-in on top of activation shaping. Evaluated on the ImageNet OOD benchmark.
[…]
When combined with Energy  Liu et al. (2020) score, these methods improve the OOD Detection performance. On Table 5 we show that applying ORA scoring after activation shaping algorithms improves the performance. Specifically combining ORA with ReAct reduces FPR95 from $25.04\%$ to $20.36\%$ highlighting both the flexibility and efficacy of our method. This demonstrates that ORA can flexibly be combined with activation shaping algorithms.
4.5 Ablation Studies
In this section, we will demonstrate the effectiveness of design choices on our score function ORA. We first justify our choice of centering in $\bm{\mu}_{\text{ID}}$ empirically, among the candidates: $\bm{\mu}_{\text{ID}}$, $\bm{\mu}_{y_{\text{pred}}}$, $\bm{\mu}_{y_{\text{target}}}$ and $\max{(\mathbf{z}_{\text{ID}}})$. Then, we compare different angle aggregation techniques across classes by replacing our $\max{\left({\{\theta{y,y^{\prime}}\}_{y^{\prime}\in\mathcal{Y},y^{\prime}\neq y}}\right)}$ with mean and min across classes.
Table 6: Ablation on different centering strategies. Evaluated on both CIFAR-10 and ImageNet OOD benchmarks.
[…]
Centering with $\mu_{\text{ID}}$ incorporates ID-statistics without biasing towards one particular class.
Table 6 shows the performance comparison between centerings with respect to different points. Using the relative angle with respect to the predicted ($\bm{\mu}_{y_{\text{pred}}}$) or target ($\bm{\mu}_{y_{\text{target}}}$) class centroid induce a bias towards the corresponding class, which in the end hinders the compatibility between angles coming across classes. On the other hand, using $\max{(\mathbf{z}_{\text{ID}}})$ shifts every feature representation to the same orthant, reducing to simply computing the absolute distance between feature representations and the decision boundaries, which is agnostic from the in-distribution feature statistics. We observe a significant improvement in performance when computing relative angles using $\mu_{\text{ID}}$, demonstrating the importance of incorporating in-distribution (ID) statistics when measuring the relationship between feature representations and decision boundaries. ORA with $\mu_{\text{ID}}$ centering improves the FPR95 by up to $1.45\%$ and $7.4\%$ on CIFAR-10 and ImageNet respectively while also improving the AUROC for both benchmarks.
[…]
Looking at the furthest class is better for ID/OOD separation.
On Table 7 we explored different ways to aggregate class specific angles. Originally, we devise our score function to return the maximum relative angle discrepancy between the feature representation across decision boundaries. Intuitively, this suggests that considering the furthest possible class that a feature belongs from the mean of in-distribution’s perspective is effective to distinguish OOD from ID. On the other hand, comparing the minimum focuses on the smallest relative angle, reducing the separability significantly. Table 7 demonstrates taking the maximum across classes clearly outperforms mean and min aggregations, improving FPR95 and AUROC metrics on both benchmarks. Specifically the difference is higher on our large-scale experiments reducing the FPR95 by $7.72\%$ and increasing the AUROC by $1.39\%$ compared to the second best aggregation.
5 Conclusion
In this paper, we introduce a novel angle-based OOD detection score function. As a post-hoc measure of model confidence, ORA offers several key advantages: it is (i) hyperparameter-free, (ii) model-agnostic and (iii) scale-invariant. These features allow ORA to be applied to arbitrary pretrained models and used in conjunction with existing activation shaping algorithms, enhancing the performance. Notably, its scale-invariant nature enables simple aggregation of multiple models’ confidences through score summation, allowing a creation of an effective model ensemble for OOD detection. Our extensive experiments demonstrate that ORA achieves state-of-the-art performance, using the relationship between the feature representations and decision boundaries relative to the ID statistics effectively. Despite the state-of-the-art performance achieved by ORA, one possible limitation might be the use of the mean alone to capture the ID statistics in our score. As a future direction, we plan to mitigate this possible limitation by incorporating multiple relative angles to better capture the ID-statistics beyond the mean, aiming to further improve OOD detection performance. We hope that our approach inspires further research into geometric interpretations of model confidence for OOD detection.
Appendix A Additional Results
A.1 OOD Detection on CIFAR-10 and ImageNet Benchmarks
@@
Table 8: ORA achieves state-of-the-art performance on CIFAR-10 OOD benchmark. Evaluated on ResNet-18 with FPR95 and AUROC. $\uparrow$ indicates that larger values are better and vice versa. Best performance highlighted in bold. Methods with * are hyperparameter-free.
@@
ORA * 3.53 99.16 8.36 98.28 23.40 94.88 8.58 98.34 10.97 97.67
Table 9: ORA achieves state-of-the-art performance on ImageNet OOD benchmark. Evaluated on ResNet-50 with FPR95 and AUROC. $\uparrow$ indicates that larger values are better and vice versa. Best performance highlighted in bold. Methods with * are hyperparameter-free.
@@
ORA * 12.27 97.42 31.80 92.85 40.71 90.10 15.39 96.68 25.04 94.26
Table 8 and Table 9 shows the performance of ORA along with the 9 baselines on CIFAR-10 and ImageNet OOD Benchmarks, respectively. All the baselines on CIFAR-10 use ResNet-18 architecture and on ImageNet use ResNet-50. Our proposed method reaches state-of-the-art performance on both benchmarks, reducing the FPR95 on average by $7.74\%$ on Imagenet and $0.88\%$ on CIFAR10. In the following, we provide a detailed analysis of these results.
@@
ORA continues to show the success of distance-based methods over logit-based methods. Logit-based scoring methods MSP  Hendrycks and Gimpel [2022], Energy  Liu et al. [2020] are one of the earliest baselines proving their success on measuring model’s confidences. MSP measures the maximum softmax probability as its score while Energy does a logsumexp operation on the logits. Recent distance-based methods like KNN+  Sun et al. [2022] and fDBD  Liu and Qin [2024] outperforms the early logit-based ones. Similarly, ORA achieves significantly better performance on both benchmarks, reducing the FPR95 up to $49.81\%$ and $41.91\%$ while improving the AUROC up to $7.53\%$ and $12.27\%$ on CIFAR-10 and ImageNet OOD benchmarks.
@@
ORA improves on the recent success of methods using contrastively learned features. Table 8 and 9 show the success of recent methods CSI  Tack et al. [2020], SSD+  Sehwag et al. [2021], KNN+  Sun et al. [2022] and fDBD  Liu and Qin [2024] that utilizes contrastively learned representations over the ones those do not use. We observe that the additional structure the supervised contrastive loss puts on the feature representations are particularly beneficial to the distance-based methods. ORA also benefits from more structured representations on the feature space, as it explores the relationship between the representation and the decision boundaries. Notably, ORA improves both of the metrics on both CIFAR-10 and ImageNet benchmarks, achieving the state-of-the-art performance.
A.2 Comparison with ReAct
Tables 10 and 11 compare the performance of ORA with the activation shaping method ReAct Sun et al. [2021]. ORA improves FPR95 by $1.11\%$ on CIFAR10 and $6.39\%$ on ImageNet. Additionally, combining ORA with ReAct further enhances performance on both benchmarks across both metrics, FPR95 and AUROC.
Table 10: ORA vs ReAct under ImageNet OOD benchmark.